Reading glasses of what power are needed for a person whose near point is 50.0 cm, so that she can read clearly the print of a book located at 25.0 cm? Assume a lens-eye distance of 2.00 cm.
Answer #1

Solution)

Given,

di = 50 cm

do = 25 cm

f = ?

Now,

(1/25) + (1/-50) = (1/f)

f = 50 cm=0.5 m

Now , Power of lens

P = (1/f)

P= (1/.5) = 2 dp

========÷÷